### 25 Mar 2026 v2

# Evolutionarily Stable Stackelberg Equilibrium

Abstract
We present a new solution concept called evolutionarily stable Stackelberg equilibrium (SESS). We study the Stackelberg evolutionary game setting in which there is a single leading player and a symmetric population of followers. The leader selects an optimal mixed strategy, anticipating that the follower population plays an evolutionarily stable strategy (ESS) in the induced subgame and may satisfy additional ecological conditions. We consider both leader-optimal and follower-optimal selection among ESSs, which arise as special cases of our framework. Prior approaches to Stackelberg evolutionary games either define the follower response via evolutionary dynamics or assume rational best-response behavior, without explicitly enforcing stability against invasion by mutations. We present algorithms for computing SESS in discrete and continuous games, and validate the latter empirically. Our model applies naturally to biological settings; for example, in cancer treatment the leader represents the physician and the followers correspond to competing cancer cell phenotypes.
I Introduction
While Nash equilibrium has emerged as the standard solution concept in game theory, it is often criticized as being too weak: often games contain multiple Nash equilibria (sometimes even infinitely many), and we want to select one that satisfies other natural properties. For example, one popular concept that refines Nash equilibrium is evolutionarily stable strategy (ESS). A mixed strategy in a two-player symmetric game is an evolutionarily stable strategy if it is robust to being overtaken by a mutation strategy. Formally, $\mathbf{x}^{*}$ is an ESS if for every mixed strategy $\mathbf{x}$ that differs from $\mathbf{x}^{*}$, there exists $\epsilon_{0}=\epsilon_{0}(\mathbf{x})>0$ such that, for all $\epsilon\in(0,\epsilon_{0})$,
$(1-\epsilon)u_{1}(\mathbf{x},\mathbf{x}^{*})+\epsilon u_{1}(\mathbf{x},\mathbf{x})<(1-\epsilon)u_{1}(\mathbf{x}^{*},\mathbf{x}^{*})+\epsilon u_{1}(\mathbf{x}^{*},\mathbf{x}).$
From a biological perspective, we can interpret $\mathbf{x}^{*}$ as a distribution among “normal” individuals within a population, and consider a mutation that makes use of strategy $\mathbf{x}$, assuming that the proportion of the mutation in the population is $\epsilon$. In an ESS, the expected payoff of the mutation is smaller than the expected payoff of a normal individual, and hence the proportion of mutations will decrease and eventually disappear over time, with the composition of the population returning to being mostly $\mathbf{x}^{*}$. An ESS is therefore a mixed strategy that is immune to being overtaken by mutations. ESS was initially proposed by mathematical biologists motivated by applications such as population dynamics (e.g., maintaining robustness to mutations within a population of humans or animals) [21, 20]. A common example game is the 2x2 game where strategies correspond to an “aggressive” Hawk or a “peaceful” Dove strategy. A paper has recently proposed a similar game in which an aggressive malignant cell competes with a passive normal cell for biological energy, which has applications to cancer eradication [6].
While Nash equilibrium is defined for general multiplayer games, ESS is traditionally defined specifically for two-player symmetric games. ESS is a refinement of Nash equilibrium. In particular, if $\mathbf{x}^{*}$ is an ESS, then $(\mathbf{x}^{*},\mathbf{x}^{*})$ (i.e., the strategy profile where both players play $\mathbf{x}^{*}$) is a (symmetric) Nash equilibrium [19]. Of course the converse is not necessarily true (not every symmetric Nash equilibrium is an ESS), or else ESS would be a trivial refinement. In fact, ESS is not guaranteed to exist in games with more than two pure strategies per player (while Nash equilibrium is guaranteed to exist in all finite games). For example, while rock-paper-scissors has a mixed strategy Nash equilibrium (which puts equal weight on all three actions), it has no ESS [19] (that work considers a version where payoffs are 1 for a victory, 0 for loss, and $\frac{2}{3}$ for a tie).
There exists a polynomial-time algorithm for computing Nash equilibrium (NE) in two-player zero-sum games, while for two-player non-zero-sum and multiplayer games computing an NE is PPAD-complete and it is widely conjectured that no efficient (polynomial-time) algorithm exists. However, several algorithms have been devised that perform well in practice. The problem of computing whether a game has an ESS was shown to be both NP-hard and CO-NP hard and also to be contained in $\Sigma^{P}_{2}$ (the class of decision problems that can be solved in nondeterministic polynomial time given access to an NP oracle) [7]. Subsequently it was shown that the exact complexity of this problem is that it is $\Sigma^{P}_{2}$-complete [5], and even more challenging for more than two players [2]. Note that this result is for determining whether an ESS exists (as discussed above there exist games which have no ESS), not for the complexity of computing an ESS in games for which one exists. Thus, computing an ESS is significantly more difficult than computing an NE. Several approaches have been proposed for computing ESS in two-player games [14, 1, 4, 3, 23].
A normal-form game consists of a finite set of players $N=\{1,\ldots,n\}$, a finite set of pure strategies $S_{i}$ for each player $i$, and a real-valued utility for each player for each strategy vector (aka strategy profile), $u_{i}:\times_{i}S_{i}\rightarrow\mathbb{R}$. In a symmetric normal-form game, all strategy spaces $S_{i}$ are equal and the utility functions satisfy the following symmetry condition: for every player $i\in N$, pure strategy profile $(s_{1},\ldots,s_{n})\in S^{n}$, and permutation $\pi$ of the players,
$u_{i}(s_{1},\ldots,s_{n})=u_{\pi(i)}(s_{\pi(1)},\ldots,s_{\pi(n)}).$
This allows us to remove the player index of the utility function and just write $u(s_{1},\ldots,s_{n}),$ where it is implied that the utility is for player 1 (we can simply permute the players to obtain the utilities of the other players). We write $u_{i}$ for notational convenience, but note that only a single utility function must be specified which applies to all players.
Let $\Sigma_{i}$ denote the set of mixed strategies of player $i$ (probability distributions over elements of $S_{i}$). If players follow mixed strategy profile $\mathbf{x}=(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)}),$ where $\mathbf{x}^{(i)}\in\Sigma_{i}$, the expected payoff to player $i$ is
$u_{i}(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)})=\sum_{s_{1},\ldots,s_{n}\in S}x^{(1)}_{s_{1}}\cdots x^{(n)}_{s_{n}}u_{i}(s_{1},\ldots,s_{n}).$
We write $u_{i}(\mathbf{x})=u_{i}(\mathbf{x}^{(i)},\mathbf{x}^{(-i)})$, where $\mathbf{x}^{(-i)}$ denotes the vector of strategies of all players except $i$. If all players follow the same mixed strategy $\mathbf{x}$, then for all players $i$:
$u_{i}(\mathbf{x})=u_{i}(\mathbf{x},\ldots,\mathbf{x})=\sum_{s_{1},\ldots,s_{n}\in S}x_{s_{1}}\cdots x_{s_{n}}u_{i}(s_{1},\ldots,s_{n}).$
Definition 1
A mixed strategy profile $\mathbf{x}^{\star}$ is a Nash equilibrium if for each player $i\in N$ and for each mixed strategy $\mathbf{x}^{(i)}\in\Sigma_{i}$: $u_{i}(\mathbf{x}^{*(i)},\mathbf{x}^{*(-i)})\geq u_{i}(\mathbf{x}^{(i)},\mathbf{x}^{*(-i)}).$
Definition 2
A mixed strategy profile $\mathbf{x}^{\star}$ in a symmetric normal-form game in a symmetric Nash equilibrium if it is a Nash equilibrium and: $\mathbf{x}^{*(1)}=\mathbf{x}^{*(2)}=\ldots=\mathbf{x}^{*(n)}.$
Definition 3
A mixed strategy $\mathbf{x}^{\star}\in\Sigma_{1}$ is evolutionarily stable in a symmetric normal-form game if for each mixed strategy $\mathbf{x}\neq\mathbf{x}^{\star}$ exactly one of the following holds:
  1. 1.
    $u_{1}(\mathbf{x}^{*},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})>u_{1}(\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})$,
  2. 2.
    $u_{1}(\mathbf{x}^{*},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})=u_{1}(\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})$ and $u_{1}(\mathbf{x}^{*},\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})>u_{1}(\mathbf{x},\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})$.
It has been proven that every symmetric normal-form game has at least one symmetric Nash equilibrium [24]. It is clear from Definition 3 that every evolutionarily stable strategy in a symmetric normal-form game must be a symmetric Nash equilibrium (SNE). Thus, a natural approach for ESS computation is to first compute SNE and then perform subsequent procedures to determine whether they are ESS.
Now we present the Stackelberg evolutionary game setting. The leader’s set of pure strategies is $\mathcal{M}=\{1,\ldots,m\}$ and the follower phenotypes are $\mathcal{P}=\{1,\ldots,n\}.$ The leader’s utility function is given by $\mathbf{A_{L}}\in\mathbb{R}^{m\times n}$, where $\mathbf{A_{L}}(\ell,i)$ gives the payoff to the leader when the leader plays pure strategy $\ell\in\mathcal{M}$ and the follower population state is phenotype $i\in\mathcal{P}.$ The leader’s expected utility when the leader plays mixed strategy $\boldsymbol{\sigma}$ and the follower plays mixed strategy $\mathbf{x}$ is $U_{L}(\boldsymbol{\sigma},\mathbf{x})=\sum_{\ell=1}^{m}\sum_{i=1}^{n}\sigma_{\ell}x_{i}\mathbf{A_{L}}(\ell,i).$ The follower payoff tensor is $\mathbf{A_{F}}\in\mathbb{R}^{m\times n\times n},$ where $\mathbf{A_{F}}(\ell,i,j)$ is the payoff (fitness) to a follower of phenotype $i\in\mathcal{P}$ when the leader plays pure strategy $\ell\in\mathcal{M}$ and the interacting follower has phenotype $j\in\mathcal{P}.$ For each fixed $\ell\in\mathcal{M},$ $\mathbf{A_{F}}(\ell,\cdot,\cdot)$ is an $n\times n$ symmetric payoff matrix. After the leader commits to mixed strategy $\boldsymbol{\sigma}\in\Delta(\mathcal{M})$, the induced follower payoff matrix is $\mathbf{B}^{\boldsymbol{\sigma}}\in\mathbb{R}^{n\times n}$ with $\mathbf{B}^{\boldsymbol{\sigma}}=\sum_{\ell=1}^{m}\sigma_{\ell}\mathbf{A_{F}}(\ell,\cdot,\cdot).$ The follower population then plays a symmetric evolutionary game with payoff matrix $\mathbf{B}^{\boldsymbol{\sigma}}.$ A population state $\mathbf{x}\in\Delta(\mathcal{P})$ is admissible for $\boldsymbol{\sigma}$ iff it is an ESS of $\mathbf{B}^{\boldsymbol{\sigma}}.$ Let $\mathcal{E}(\boldsymbol{\sigma})$ denote the set of ESSs of $\mathbf{B}^{\boldsymbol{\sigma}},$ and let $g^{\boldsymbol{\sigma}}:\Delta(\mathcal{P})\rightarrow\mathbb{R}$ denote the follower’s ESS selection function. Finally let $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}g^{\boldsymbol{\sigma}}(\mathbf{x}).$ The leader maximizes utility anticipating that the follower outcome is selected from the correspondence $\mathcal{G}(\boldsymbol{\sigma}).$
We now present our general definition for evolutionarily stable Stackelberg equilibrium (SESS) in Definition 4. We also define two natural special cases. In the first case, the followers play the ESS that is most beneficial to the leader. We refer to this as an optimistic evolutionarily stable Stackelberg equilibrium (OSESS). In this case we have $g^{\boldsymbol{\sigma}}(\mathbf{x})=U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ We also consider pessimistic evolutionarily stable Stackelberg equilibrium (PSESS), in which the followers play the ESS that is worst for the leader. While the followers are not rationally trying to harm the leader, we can view PSESS as SESS that is robust to worst-case evolutionary stability. For PSESS we have $g^{\boldsymbol{\sigma}}(\mathbf{x})=-U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ OSESS and PSESS are defined below in Definitions 56. For OSESS we have $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})$, and for PSESS $\mathcal{G}(\boldsymbol{\sigma})=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x}).$ It is clear that OSESS and PSESS are refinements of SESS (since they specify a particular function $g^{\boldsymbol{\sigma}}(\mathbf{x})$). SESS can be viewed as a refinement of Stackelberg equilibrium in which the follower response correspondence is restricted to evolutionarily stable strategies in the induced population game (Proposition 1).
Definition 4 (Evolutionarily Stable Stackelberg Equilibrium)
Given a selection correspondence $\mathcal{G}(\boldsymbol{\sigma})\subseteq\mathcal{E}(\boldsymbol{\sigma})$, $(\boldsymbol{\sigma}^{*},\mathbf{x}^{*})$ is an evolutionarily stable Stackelberg equilibrium if
$\boldsymbol{\sigma}^{*}\in\operatorname*{arg\,max}_{\boldsymbol{\sigma}\in\Delta(\mathcal{M})}\;\max_{\mathbf{x}\in\mathcal{G}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})\quad\text{and}\quad\mathbf{x}^{*}\in\mathcal{G}(\boldsymbol{\sigma}^{*}).$
Definition 5 (Optimistic SESS)
A pair $(\boldsymbol{\sigma}^{*},\mathbf{x}^{*})\in\Delta(\mathcal{M})\times\Delta(\mathcal{P})$ is an optimistic evolutionarily stable Stackelberg equilibrium if
$(\boldsymbol{\sigma}^{*},\mathbf{x}^{*})\in\operatorname*{arg\,max}_{\boldsymbol{\sigma}\in\Delta(\mathcal{M}),\;\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x}).$
Definition 6 (Pessimistic SESS)
A pair $(\boldsymbol{\sigma}^{*},\mathbf{x}^{*})\in\Delta(\mathcal{M})\times\Delta(\mathcal{P})$ is a pessimistic evolutionarily stable Stackelberg equilibrium if
$\boldsymbol{\sigma}^{*}\in\operatorname*{arg\,max}_{\boldsymbol{\sigma}\in\Delta(\mathcal{M})}\;\min_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\mathbf{x})\text{ and }\mathbf{x}^{*}\in\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{E}(\boldsymbol{\sigma}^{*})}U_{L}(\boldsymbol{\sigma}^{*},\mathbf{x}).$
Proposition 1
Every evolutionarily stable Stackelberg equilibrium $(\boldsymbol{\sigma}^{*},\boldsymbol{x}^{*})$ is a Stackelberg equilibrium.
Proof:
Suppose that $(\boldsymbol{\sigma}^{*},\boldsymbol{x}^{*})$ is an evolutionarily stable Stackelberg equilibrium. By Definition 4,
$\boldsymbol{\sigma}^{*}\in\arg\max_{\boldsymbol{\sigma}\in\Delta(M)}\max_{\boldsymbol{x}\in\mathcal{G}(\boldsymbol{\sigma})}U_{L}(\boldsymbol{\sigma},\boldsymbol{x})\quad\text{and}\quad\boldsymbol{x}^{*}\in\mathcal{G}(\boldsymbol{\sigma}^{*}).$
Since $\mathcal{G}(\boldsymbol{\sigma}^{*})\subseteq\mathcal{E}(\boldsymbol{\sigma}^{*})$, it follows that
$\boldsymbol{x}^{*}\in\mathcal{E}(\boldsymbol{\sigma}^{*}),$
so the follower outcome $\boldsymbol{x}^{*}$ is an evolutionarily stable strategy in the induced follower game under $\boldsymbol{\sigma}^{*}$.
Moreover, $\boldsymbol{\sigma}^{*}$ maximizes the leader’s utility over all strategies, anticipating that the follower response is selected from $\mathcal{G}(\boldsymbol{\sigma})$. Since $\boldsymbol{x}^{*}\in\mathcal{G}(\boldsymbol{\sigma}^{*})$, the pair $(\boldsymbol{\sigma}^{*},\boldsymbol{x}^{*})$ satisfies the Stackelberg optimality and consistency conditions. Hence it is a Stackelberg equilibrium. ∎
II Continuous Stackelberg–ESS
We now extend the Stackelberg–ESS framework to continuous Stackelberg evolutionary games, which arise naturally in biological and control applications such as cancer treatment. In this setting, the leader selects a continuous decision variable, while the followers correspond to an evolving population whose response is described by coupled ecological and evolutionary dynamics.
Let $\mathcal{M}\subseteq\mathbb{R}_{+}^{d}$ denote the leader decision set (e.g., treatment doses or control parameters). For a given leader decision $\mathbf{m}\in\mathcal{M}$, the followers are described by a collection of continuous state variables, including population sizes and traits. We denote the follower outcome by $\mathbf{z}=(\mathbf{x},\mathbf{u})\in\mathcal{Z}$, where $\mathbf{x}\in\mathbb{R}_{+}^{n}$ represents population abundances and $\mathbf{u}\in\mathcal{U}\subseteq\mathbb{R}^{n}$ represents trait or strategy variables. As prior work has done we will assume that $\mathcal{U}=[0,1]^{n}.$ In the context of cancer treatment $d$ denotes the number of drugs and $n$ denotes the number of cancer cell phenotypes. The leader objective is given by a function $Q:\mathcal{M}\times\mathcal{Z}\to\mathbb{R}$. For each $\mathbf{m}\in\mathcal{M}$, the leader induces an eco–evolutionary system governing the dynamics of $(\mathbf{x},\mathbf{u})$.
For a fixed leader decision $\mathbf{m}$, we define the set of admissible follower responses $\mathcal{E}(\mathbf{m})\subseteq\mathcal{Z}$ as the set of eco–evolutionarily stable outcomes under $\mathbf{m}$. Informally, a follower outcome $\mathbf{z}^{*}=(\mathbf{x}^{*},\mathbf{u}^{*})$ is admissible if it satisfies any required ecological equilibrium conditions and is stable against invasion by rare mutant phenotypes in the induced eco–evolutionary system. When multiple evolutionarily stable outcomes exist for the same leader decision, we model their selection via a correspondence $\mathcal{G}(\mathbf{m})\subseteq\mathcal{E}(\mathbf{m})$. We assume that $\mathcal{E}(\boldsymbol{m})\neq\emptyset$ for all $\boldsymbol{m}\in M$, so that the follower response correspondence is well-defined.
We now define continuous analogues of Stackelberg–ESS. It is again clear that in the continuous setting OSESS and PSESS are refinements of SESS, and SESS is a refinement of Stackelberg equilibrium in which follower responses are restricted to evolutionarily stable strategies.
Definition 7 (Continuous Stackelberg–ESS)
Let $\mathcal{G}(\mathbf{m})\subseteq\mathcal{E}(\mathbf{m})$ be a selection correspondence. A pair $(\mathbf{m}^{*},\mathbf{z}^{*})\in\mathcal{M}\times\mathcal{Z}$ is a continuous Stackelberg–ESS if
$\mathbf{m}^{*}\in\operatorname*{arg\,max}_{\mathbf{m}\in\mathcal{M}}\ \max_{\mathbf{z}\in\mathcal{G}(\mathbf{m})}Q(\mathbf{m},\mathbf{z})\quad\text{and}\quad\mathbf{z}^{*}\in\mathcal{G}(\mathbf{m}^{*}).$
Definition 8 (Optimistic continuous Stackelberg–ESS)
A pair $(\mathbf{m}^{*},\mathbf{z}^{*})\in\mathcal{M}\times\mathcal{Z}$ is an optimistic continuous Stackelberg–ESS if
$(\mathbf{m}^{*},\mathbf{z}^{*})\in\operatorname*{arg\,max}_{\mathbf{m}\in\mathcal{M},\ \mathbf{z}\in\mathcal{E}(\mathbf{m})}Q(\mathbf{m},\mathbf{z}).$
Definition 9 (Pessimistic continuous Stackelberg–ESS)
A pair $(\mathbf{m}^{*},\mathbf{z}^{*})\in\mathcal{M}\times\mathcal{Z}$ is a pessimistic continuous Stackelberg–ESS if
$\mathbf{m}^{*}\in\operatorname*{arg\,max}_{\mathbf{m}\in\mathcal{M}}\ \min_{\mathbf{z}\in\mathcal{E}(\mathbf{m})}Q(\mathbf{m},\mathbf{z})\text{ and }\mathbf{z}^{*}\in\operatorname*{arg\,min}_{\mathbf{z}\in\mathcal{E}(\mathbf{m}^{*})}Q(\mathbf{m}^{*},\mathbf{z}).$
Proposition 2
Every continuous evolutionarily stable Stackelberg equilibrium is a Stackelberg equilibrium.
Proof:
Suppose that $(\boldsymbol{m}^{*},\boldsymbol{z}^{*})$ is a continuous evolutionarily stable Stackelberg equilibrium. By Definition 7,
$\boldsymbol{m}^{*}\in\arg\max_{\boldsymbol{m}\in M}\max_{\boldsymbol{z}\in\mathcal{G}(\boldsymbol{m})}Q(\boldsymbol{m},\boldsymbol{z})\quad\text{and}\quad\boldsymbol{z}^{*}\in\mathcal{G}(\boldsymbol{m}^{*}).$
Since $\mathcal{G}(\boldsymbol{m}^{*})\subseteq\mathcal{E}(\boldsymbol{m}^{*})$, it follows that $\boldsymbol{z}^{*}\in\mathcal{E}(\boldsymbol{m}^{*}),$ so the follower outcome $\boldsymbol{z}^{*}$ is eco–evolutionarily stable under $\boldsymbol{m}^{*}$. Moreover, $\boldsymbol{m}^{*}$ maximizes the leader’s objective over all decisions, anticipating that the follower response is selected from $\mathcal{G}(\boldsymbol{m})$. Since $\boldsymbol{z}^{*}\in\mathcal{G}(\boldsymbol{m}^{*})$, the pair $(\boldsymbol{m}^{*},\boldsymbol{z}^{*})$ satisfies the Stackelberg optimality and consistency conditions. Hence it is a Stackelberg equilibrium. ∎
We now present the Stackelberg evolutionary game formulation of cancer treatment from prior work [18]. In this formulation (Equation 1), the objective Q corresponds to the quality of life function that is maximized by the leader. For each cell type $i$, there is a fitness function $G_{i}(u_{i},\mathbf{m},\mathbf{x})$ that the follower is trying to maximize. We assume that the dynamics of the population $\mathbf{x}$ are governed by $\dot{\mathbf{x}}=\mathbf{G}(t)\mathbf{x}.$ In order to ensure that we are in equilibrium of the ecological dynamics we must have that $\dot{x_{i}}=0$ for all $i.$ Note that in this formulation the follower is playing a best response by maximizing its fitness function. By contrast, in Stackelberg–ESS we assume that the behavior of the cancer cells is determined by evolutionary stability conditions rather than explicit optimization. An algorithm for solving this problem has been presented that is based on a nonconvex quadratic program formulation [8]. This algorithm has been demonstrated to quickly compute a global optimal solution on a realistic example problem proposed by prior work [18].
$\max_{\mathbf{m}^{*},\mathbf{u}^{*},\mathbf{x}^{*}}$ $Q(\mathbf{m}^{*},\mathbf{u}^{*},\mathbf{x}^{*})$ (1)
s.t. $\dot{x}_{i}^{*}=0,\quad i=1,\ldots,n,$
$u_{i}^{*}\in\arg\max_{u_{i}\in[0,1]}G_{i}(u_{i},\mathbf{m}^{*},\mathbf{x}^{*}),\quad i=1,\ldots,n,$
$\mathbf{m}^{*}\geq\mathbf{0},\quad\mathbf{x}^{*}\geq\mathbf{0},\quad\mathbf{0}\leq\mathbf{u}^{*}\leq\mathbf{1}.$
The formulation for optimistic continuous Stackelberg-ESS is given by Equation 2. This formulation is similar to the previous one for standard Stackelberg equilibrium except that now instead of maximizing the fitness functions the follower ensures that their strategy is an ESS in the induced subgame given the leader is playing $m^{*}.$ We will focus our attention on the problem of computing OSESS under this formulation, since it is perhaps the easiest case of SESS to solve since both players are aligned in maximizing $Q.$
$\max_{\mathbf{m}^{*},\mathbf{u}^{*},\mathbf{x}^{*}}$ $Q(\mathbf{m}^{*},\mathbf{u}^{*},\mathbf{x}^{*})$ (2)
s.t. $\dot{x}_{i}^{*}=0,\quad i=1,\ldots,n,$
$(\mathbf{x}^{*},\mathbf{u}^{*})\in\mathcal{E}(\mathbf{m}^{*}),$
$\mathbf{m}^{*}\geq\mathbf{0},\quad\mathbf{x}^{*}\geq\mathbf{0},\quad\mathbf{0}\leq\mathbf{u}^{*}\leq\mathbf{1}.$
Equation 3 provides an equivalent reformulation of Equation 2. Note that we use the continuous-trait definition of ESS for our setting [16, 25, 11] as opposed to the discrete ESS definition presented previously. Given fixed $\mathbf{m}$, a resident state $(\mathbf{x}^{*},\mathbf{u}^{*})$ is an ESS iff $\mathbf{x}^{*}$ is an ecological equilibrium (i.e., $\dot{\mathbf{x}}^{*}=\mathbf{0}$) and no rare mutant can grow, i.e., $\sup_{u_{i}}G_{i}(u_{i},\mathbf{m},\mathbf{x}^{*})\leq 0\mbox{ }\forall i.$ Recall that $\dot{\mathbf{x}}=\mathbf{G}\mathbf{x}.$ So the constraint $\dot{x}^{*}_{i}=0$ is equivalent to $G_{i}(\mathbf{u}^{*}_{i},\mathbf{m}^{*},\mathbf{x}^{*})x^{*}_{i}=0.$ Recall that $(\mathbf{x}^{*},\mathbf{u}^{*})\in\mathcal{E}(\mathbf{m})$ iff $(\mathbf{x}^{*},\mathbf{u}^{*})$ is evolutionarily stable in the eco-evolutionary system induced by $\mathbf{m}.$ This is true iff for all $i$ and $u_{i}$ a mutation to $u_{i}$ does not lead to positive fitness (growth rate), i.e., $G_{i}(u_{i},\mathbf{m},\mathbf{x}^{*})\leq 0.$
$\max_{\mathbf{m}^{*},\mathbf{u}^{*},\mathbf{x}^{*}}$ $Q(\mathbf{m}^{*},\mathbf{u}^{*},\mathbf{x}^{*})$ (3)
s.t. $x_{i}^{*}\,G_{i}(u_{i}^{*},\mathbf{m}^{*},\mathbf{x}^{*})=0,\quad i=1,\dots,n,$
$G_{i}(u_{i},\mathbf{m}^{*},\mathbf{x}^{*})\leq 0\quad\forall u_{i}\in[0,1],\;i=1,\dots,n,$
$\mathbf{m}^{*}\geq\mathbf{0},\quad\mathbf{x}^{*}\geq\mathbf{0},\quad\mathbf{0}\leq\mathbf{u}^{*}\leq\mathbf{1}.$
III Algorithms
We first present an algorithm for computing OSESS in normal-form games, followed by an algorithm for continuous-trait games. Pseudocode for the main algorithm in the normal-form setting is given in Algorithm 1. We compute an OSESS by enumerating candidate follower supports and for each support computing a Stackelberg equilibrium of the game where the follower is restricted to strategies with the given support. This is done by using the subroutine given in Algorithm 2. Given the Stackelberg equilibrium $(\boldsymbol{\sigma},\boldsymbol{x}),$ we next check whether $\boldsymbol{x}$ is an ESS in the subgame induced by the leader’s strategy $\boldsymbol{\sigma}.$ This is done in Algorithm 3. If the strategies do constitute an ESS, then we have found an evolutionarily stable Stackelberg equilibrium (SESS). We then calculate the objective $\boldsymbol{\sigma}^{\top}\mathbf{A}_{L},$ and return the SESS with highest value as an OSESS. The algorithm can be easily adapted to output the first SESS found, all SESSs found, or an SESS that optimizes a different objective. The values used for all numerical tolerance parameters are given in Table I.
Algorithm 2 computes a Stackelberg equilibrium strategy with given support $T$ by solving a quadratically-constrained quadratic program (QCQP). The constraints $\boldsymbol{x}_{i}\geq\epsilon_{s}$ ensure that the pure strategies in the support are played with positive probability. The objective ensures that the leader’s payoff is maximized, and the final two sets of constraints ensure that the follower is best-responding to the leader’s strategy. Our algorithm is implicitly assuming that the game is nondegenerate and that there exists at least one Stackelberg equilibrium for each follower support. This is a common assumption in equilibrium-finding algorithms [22, 12, 15], and has been studied recently in the context of ESS computation [10, 9]. If the game is degenerate with an infinite continuum of equilibria our algorithm will fail to find all of them, and therefore may fail to find all SESSs; however it may still find one. Algorithm 3 also involves solving a QCQP. In practice we will be solving nonconvex QCQPs with Gurobi’s solver, which can solve small problem instances quickly despite their computational complexity. The QCQP checks whether any mutant strategy $\mathbf{y}$ can successfully invade against a population following strategy $\mathbf{x}$ in the game induced by the leader’s strategy $\mathbf{\sigma}.$ The parameter $\epsilon_{p}$ allows for a small amount of numerical imprecision, and the final constraint ensures that $\mathbf{y}$ is numerically distinct from $\mathbf{x}$ by using numerical separation parameter $\delta.$
Algorithm 1 ComputeDiscreteOSESS
0: Follower payoff tensor $\mathbf{A}_{F}\in\mathbb{R}^{m\times n\times n}$
0: Leader payoff matrix $\mathbf{A}_{L}\in\mathbb{R}^{m\times n}$
0: Minimum support mass $\epsilon_{s}$, ESS tolerances $\epsilon_{p},\delta$
0: OSESS $(\boldsymbol{\sigma}^{\star},\boldsymbol{x}^{\star})$
[…]
3:for all nonempty supports $T\subseteq P$ do
4:  $(status,\boldsymbol{\sigma},\boldsymbol{x})\leftarrow\textsc{SolveSESupport}(\mathbf{A}_{F},\mathbf{A}_{L},T,\epsilon_{s})$
5:  if $status=\texttt{FEASIBLE}$ then
6:   if $\textsc{IsESS}(\mathbf{A}_{F},\boldsymbol{\sigma},\boldsymbol{x},\epsilon_{p},\delta)$ then
7:    $v\leftarrow\boldsymbol{\sigma}^{\top}\mathbf{A}_{L}\boldsymbol{x}$
8:    if $v>v^{\star}$ then
9:     $v^{\star}\leftarrow v$
10:     $(\boldsymbol{\sigma}^{\star},\boldsymbol{x}^{\star})\leftarrow(\boldsymbol{\sigma},\boldsymbol{x})$
11:    end if
12:   end if
13:  end if
14:end for
15:return $(\boldsymbol{\sigma}^{\star},\boldsymbol{x}^{\star})$
Algorithm 2 SolveSESupport
0: Follower payoff tensor $\mathbf{A}_{F}\in\mathbb{R}^{m\times n\times n}$
0: Leader payoff matrix $\mathbf{A}_{L}\in\mathbb{R}^{m\times n}$
0: Support $T\subseteq P$, minimum support mass $\epsilon_{s}$
[…]
1: Solve the following optimization problem:
$\max_{\boldsymbol{\sigma},\boldsymbol{x},v}$ $\boldsymbol{\sigma}^{\top}\mathbf{A}_{L}\boldsymbol{x}$
s.t. $\boldsymbol{\sigma}_{\ell}\geq 0,\quad\forall\ell\in M,$
$\sum_{\ell\in M}\boldsymbol{\sigma}_{\ell}=1,$
$\boldsymbol{x}_{i}\geq\epsilon_{s},\quad\forall i\in T,$
$\boldsymbol{x}_{j}=0,\quad\forall j\notin T,$
$\sum_{i\in P}\boldsymbol{x}_{i}=1,$
$\sum_{\ell\in M}\sum_{k\in P}\boldsymbol{\sigma}_{\ell}\mathbf{A}_{F}(\ell,i,k)\boldsymbol{x}_{k}=v,\quad\forall i\in T,$
$\sum_{\ell\in M}\sum_{k\in P}\boldsymbol{\sigma}_{\ell}\mathbf{A}_{F}(\ell,j,k)\boldsymbol{x}_{k}\leq v,\quad\forall j\notin T.$
2:if the problem is feasible then
3:  return $(\texttt{FEASIBLE},\boldsymbol{\sigma},\boldsymbol{x})$
4:else
5:  return $(\texttt{INFEASIBLE},\varnothing,\varnothing)$
6:end if
Algorithm 3 IsESS
0: Follower payoff tensor $\mathbf{A}_{F}\in\mathbb{R}^{m\times n\times n}$
0: Leader mixed strategy $\boldsymbol{\sigma}\in\Delta(M)$
0: Follower mixed strategy $\boldsymbol{x}\in\Delta(P)$
0: Tolerances $\epsilon_{p},\delta$
0:TRUE if $\boldsymbol{x}$ is an ESS of the induced follower game; otherwise FALSE
1: Define the induced follower payoff matrix $\mathbf{B}\in\mathbb{R}^{n\times n}$:
$\mathbf{B}_{ij}\leftarrow\sum_{\ell\in M}\boldsymbol{\sigma}_{\ell}\mathbf{A}_{F}(\ell,i,j),\qquad\forall i,j\in P.$
[…]
3: Solve the following optimization problem:
$\max_{\boldsymbol{y}}$ $\boldsymbol{y}^{\top}\mathbf{B}\boldsymbol{y}-\boldsymbol{x}^{\top}\mathbf{B}\boldsymbol{y}$
s.t. $\boldsymbol{y}_{j}\geq 0,\quad\forall j\in P,$
$\sum_{j\in P}\boldsymbol{y}_{j}=1,$
$\left|\boldsymbol{y}^{\top}\mathbf{B}\boldsymbol{x}-v\right|\leq\epsilon_{p},$
$\|\boldsymbol{y}-\boldsymbol{x}\|^{2}_{2}\geq\delta.$
4:if the optimal objective value $\leq\epsilon_{p}$ then
5:  return TRUE
6:else
7:  return FALSE
8:end if
Pseudocode for the main algorithm in continuous setting is given in Algorithm 4. The algorithm follows a generate-and-certify structure: first a candidate solution satisfying the KKT conditions of the invasion maximization problems is computed, and then a global optimization check verifies that no mutant phenotype can invade. The main subroutine Algorithm 5 generates a candidate OSESS by solving the optimization problem presented in Equation 3. The formulation finds a solution that satisfies the KKT optimality conditions, though it is not guaranteed to be a globally optimal solution. To test if this candidate solution is globally optimal we perform the ex-post certification procedure described in Algorithm 6, using a numerical tolerance of $\epsilon_{inv}.$ We set $\epsilon_{inv}=10^{-3},$ which is biologically negligible but significantly larger than Gurobi’s default tolerance of $10^{-6}.$ Note that if the quality of life functions $G_{i}$ are concave in $u_{i}$ then the KKT conditions are also sufficient. In our experiments we will use previously-studied $Q$ and $G_{i}$ functions, and the algorithm will involve solving nonconvex QCQPs.
Algorithm 4 ComputeContinuousOSESS
0: Leader objective $Q(\mathbf{m},\mathbf{u},\mathbf{x})$
0: Fitness functions $\{G_{i}(u_{i},\mathbf{m},\mathbf{x})\}_{i=1}^{n}$
0: Feasible set $\mathcal{M}$, tolerance $\epsilon_{\mathrm{inv}}$
0: OSESS $(\mathbf{m}^{\star},\mathbf{u}^{\star},\mathbf{x}^{\star})$ if one is found
[…]
2:for $i=1$ to $n$ do
3:  $G_{i}^{\max}\leftarrow\textsc{Certify}(i,\mathbf{m}^{\star},\mathbf{x}^{\star})$
4:end for
5:if $\max_{i}G_{i}^{\max}\leq\epsilon_{\text{inv}}$ then
6:  return $(\mathbf{m}^{\star},\mathbf{u}^{\star},\mathbf{x}^{\star})$
7:else
8:  return FAILURE
9:end if
Algorithm 5 GenOSESS
0: Leader objective $Q(\mathbf{m},\mathbf{u},\mathbf{x})$
0: Fitness functions $\{G_{i}(u_{i},\mathbf{m},\mathbf{x})\}_{i=1}^{n}$
0: Feasible set $\mathcal{M}$
0: Candidate solution $(\mathbf{m}^{\star},\mathbf{u}^{\star},\mathbf{x}^{\star})$
1: Solve the following optimization problem:
$\max_{\mathbf{m},\mathbf{u},\mathbf{x},\mathbf{u}^{\prime},\boldsymbol{\lambda}^{L},\boldsymbol{\lambda}^{U}}$ $Q(\mathbf{m},\mathbf{u},\mathbf{x})$
s.t. $x_{i}G_{i}(u_{i},\mathbf{m},\mathbf{x})=0,\qquad i=1,\ldots,n,$
[…]
$G_{i}(u^{\prime}_{i},\mathbf{m},\mathbf{x})\leq 0,\qquad i=1,\ldots,n,$
$\frac{\partial G_{i}}{\partial u_{i}}(u^{\prime}_{i},\mathbf{m},\mathbf{x})+\lambda_{i}^{L}-\lambda_{i}^{U}=0,\ i=1,\ldots,n,$
$\lambda_{i}^{L}u^{\prime}_{i}=0,\qquad i=1,\ldots,n,$
$\lambda_{i}^{U}(1-u^{\prime}_{i})=0,\qquad i=1,\ldots,n,$
$\lambda_{i}^{L}\geq 0,\;\lambda_{i}^{U}\geq 0,\qquad i=1,\ldots,n.$
2:return optimal solution $(\mathbf{m}^{\star},\mathbf{u}^{\star},\mathbf{x}^{\star})$
Algorithm 6 Certify
0: Phenotype index $i\in\{1,\ldots,n\}$
0: Fixed $(\mathbf{m}^{\star},\mathbf{x}^{\star})$
[…]
1: Computing globally optimal solution:
$\max_{u_{i}}$ $G_{i}(u_{i},\mathbf{m}^{\star},\mathbf{x}^{\star})$
s.t. $0\leq u_{i}\leq 1$
2:return optimal objective $G_{i}^{\max}$
Parameter Value
$\epsilon_{s}$ $10^{-4}$
$\epsilon_{p}$ $10^{-5}$
$\delta$ $10^{-2}$
$\epsilon_{\mathrm{inv}}$ $10^{-3}$
TABLE I: Tolerance parameter values for the algorithms.
IV Experiments
We experiment with our continuous algorithm on an evolutionary cancer game model previously studied [18]. Note that prior work for this model has computed a Stackelberg equilibrium (SE), where the leader commits to a strategy that optimizes the quality of life function $Q$ while the cancer cell phenotypes optimize their respective fitness functions $G_{i}$ (subject to the equilibrium conditions of the ecological dynamics being satisfied) [18, 8]. In contrast, we compute an OSESS in this game, which guarantees that the follower cancer cells are following an ESS in the game induced by the leader’s strategy and no mutant trait can have positive growth (while SE does not preclude existence of mutants with positive growth). This is fundamentally a different solution concept, so the players’ payoffs may increase or decrease. The model is instantiated by the following functional forms for the fitness functions $G_{i}$ and quality of life function $Q$. Note that the model presentation is slightly different between the paper [18] and the implementation in the code repository [17], and we will use the model from the code.
$G_{0}=r_{\mathrm{max}}\left(1-\frac{\alpha_{00}x_{0}+\alpha_{01}x_{1}+\alpha_{02}x_{2}}{K}\right)-d-\frac{m_{1}}{k_{1}}-\frac{m_{2}}{k_{2}}$
$G_{1}=r_{\mathrm{max}}e^{-g_{1}u_{1}}\left(1-\frac{\alpha_{10}x_{0}+\alpha_{11}x_{1}+\alpha_{12}x_{2}}{K}\right)-d-\frac{m_{1}}{b_{1}u_{1}+k_{1}}-\frac{m_{2}}{k_{2}}$
$G_{2}=r_{\mathrm{max}}e^{-g_{2}u_{2}}\left(1-\frac{\alpha_{20}x_{0}+\alpha_{21}x_{1}+\alpha_{22}x_{2}}{K}\right)-d-\frac{m_{1}}{k_{1}}-\frac{m_{2}}{b_{2}u_{2}+k_{2}}$
$Q=Q^{\mathrm{max}}-c\left(\frac{x_{0}+x_{1}+x_{2}}{K}\right)^{2}-w_{1}m^{2}_{1}-w_{2}m^{2}_{2}-r_{1}u^{2}_{1}-r_{2}u^{2}_{2}$
The model has several parameters, whose interpretations are summarized in Table II. Note that in the code additional parameters $a_{0},a_{1},a_{2},a_{3}$ are defined, with
$\begin{bmatrix}\alpha_{00}&\alpha_{01}&\alpha_{02}\\ \alpha_{10}&\alpha_{11}&\alpha_{12}\\ \alpha_{20}&\alpha_{21}&\alpha_{22}\\ \end{bmatrix}=\begin{bmatrix}a_{0}&a_{1}&a_{1}\\ a_{2}&a_{0}&a_{3}\\ a_{2}&a_{3}&a_{0}\\ \end{bmatrix}$
Parameter Interpretation
$r_{\mathrm{max}}$ Max cell growth rate
$g_{i}$ Cost of resistance strategy (cell type) $i$
$\alpha_{ij}$ Interaction coefficient between cell types $i$ and $j$
$K$ Carrying capacity
$d$ Natural death rate
$k_{i}$ Innate resistance that may be present before drug exposure
$b_{i}$ Benefit of the evolved resistance trait in reducing therapy efficacy
$Q^{\mathrm{max}}$ Quality of life of a healthy patient
$w_{i}$ Toxicity of drug $i$
$r_{i}$ Effect of resistance rate of cell type $i$
$c$ Weight for impact of tumor burden vs. drug toxicity/treatment-induced resistance rate
TABLE II: Interpretations of model parameters
We ran experiments comparing the algorithm for computing a Stackelberg equilibrium [8] to our new algorithm for computing OSESS on a problem instance using the same parameter values that have been previously used [17], which are provided in Table III. Both approaches involve solving QCQPs using Gurobi’s nonconvex mixed-integer quadratically-constrained programming (MIQCP) solver version 11.0.3 build v11.0.3rc0 [13] with Java version 14.0.2. Gurobi’s nonconvex MIQCP solver guarantees global optimality (up to a numerical tolerance). We used Gurobi’s default numerical tolerance value of $10^{-6}.$ We used Gurobi’s default values for all settings other than DualReductions which we set to 0 since we observed improved performance by disabling aggressive presolve reductions. We used an Intel Core i7-1065G7 processor with base clock speed of 1.30 GHz (and maximum turbo boost speed of up to 3.9 GHz) with 16 GB of RAM under 64-bit Windows 11 (8 logical cores/threads).
Parameter Value
$r_{\mathrm{max}}$ 0.45
$g_{1}$ 0.5
$g_{2}$ 0.5
$a_{0}$ 1
$a_{1}$ 0.15
$a_{2}$ 0.9
$a_{3}$ 0.9
$K$ 10,000
$d$ 0.01
$k_{1}$ 5
$k_{2}$ 5
$b_{1}$ 10
$b_{2}$ 10
$Q^{\mathrm{max}}$ 1
$w_{1}$ 0.5
$w_{2}$ 0.2
$r_{1}$ 0.4
$r_{2}$ 0.4
$c$ 0.5
TABLE III: Parameter values used in experiments
The main results from our experiments are depicted in Table IV. Note that the optimal physician strategies ($\mathbf{m}^{*}$) vary slightly between the two solutions, and the optimal quality of life $Q^{*}$ is slightly higher for the OSESS solution than the SE solution. The SE algorithm ran significantly faster than the OSESS algorithm (around 2 seconds versus 2.2 minutes). Both algorithms are based on KKT necessary optimality conditions and require the application of ex-post certification procedures to ensure global optimality. For both algorithms we recompute the values of $Q$ and $G_{i}$ exactly by calculating them directly from the solutions output from Gurobi (since the values of $Q$ and $G_{i}$ in the optimizations themselves contain numerical error). For the OSESS algorithm the certification procedure obtains $G^{max}_{1}=$7.85\text{\times}{10}^{-5}$\$, $G^{max}_{2}=$1.41\text{\times}{10}^{-4}$\$, which are both significantly below $\epsilon_{inv}=10^{-3}$, so we conclude that our solution is in fact an OSESS. As it turns out, running the same certification procedure on the Stackelberg equilibrium solution obtains $G^{max}_{1}=$8.00\text{\times}{10}^{-5}$\$, $G^{max}_{2}=$1.42\text{\times}{10}^{-4}$\$, which are also both significantly below $\epsilon_{inv}=10^{-3}$. So for these particular game parameters it turns out that the SE solution obtained is also an OSESS, though this will not be the case in general.
Stackelberg Equilibrium OSESS
$m^{*}_{1}$ 0.4105 0.4003
$m^{*}_{2}$ 0.4680 0.4571
$u^{*}_{1}$ 0.2139 0.1827
$u^{*}_{2}$ 0.2856 0.2828
$x^{*}_{0}$ 5731.0481 5823.7239
$x^{*}_{1}$ 0.0087 9.5179
$x^{*}_{2}$ 950.7623 946.4278
$Q^{*}$ 0.5978 0.6029
Runtime (seconds) 2.2070 134.6910
TABLE IV: Experimental results for evolutionary cancer game
V Conclusion
Stackelberg evolutionary games have emerged as a powerful model for studying many types of biological interactions, including cancer therapy, fishery management, and pest control [18]. In these games, a rational human leader first selects their strategy, then evolutionary followers select their strategies conditional on the leader’s strategy (possibly subject to ecological population equilibrium constraints). Prior work has assumed that the follower players act to maximize fitness functions $G_{i}$. In the resulting Stackelberg equilibrium, the leader and followers are simultaneously maximizing their respective objective functions. However, this solution does not guarantee that the follower strategy is resistant to invasion by rare mutants. To ensure evolutionary stability we must impose the condition that the fitness functions are nonpositive (subject to numerical tolerance) for all possible mutations so that their populations do not continue to grow. We introduced the new solution concept evolutionarily stable Stackelberg equilibrium (SESS) in which the leader is a rational utility maximizer, while the followers employ an evolutionarily stable strategy (ESS) that is resistant to invasion. In an SESS, the leader commits to a strategy that maximizes their objective assuming the followers will play an ESS in the subgame induced by the leader’s strategy. This imposes a stricter stability requirement on the follower strategy than standard Stackelberg equilibrium, which only requires best-response behavior without ensuring evolutionary stability.
We define SESS both for discrete normal-form games, as well continuous-trait games, which arise naturally in biological and control applications such as cancer treatment. In our general formulation, the followers play an ESS in the subgame induced by the leader’s strategy that maximizes an arbitrary objective function. We also consider the special cases where the followers play the ESS that is most beneficial to the leader (OSESS) and the ESS that is worst for the leader (PSESS). OSESS is perhaps the easiest case of SESS to compute since in a sense all players are aligned in maximizing the leader’s payoff (subject to the followers playing an ESS), which simplifies the optimization to a single outer maximization problem. We present algorithms for computing OSESS in both the discrete and continuous settings. If we view the follower strategies as population frequencies, then our model is applicable to settings with a large number of follower players; e.g., there may be a large number of cancer cells where each cell has one of a small number of possible phenotypes.
We implemented our OSESS algorithm on an evolutionary cancer game model that has been previously studied and demonstrated that its runtime is reasonable in practice. It is not surprising that OSESS computation takes longer than SE computation, since OSESS is essentially SE with additional stability constraints imposed and therefore involves solving a more challenging optimization problem. As it turns out, the Stackelberg equilibrium solution for this game ended up being an OSESS as well, though this is not guaranteed in general. This suggests that for larger games perhaps a good approach would be to first compute an SE, which is more tractable than SESS, and test whether it also satisfies evolutionary stability; if not then we can continue on to the more computationally expensive SESS algorithm.